OpenStudy (ttop0816):

HELP!! 2sin^2 x - cos^2 x - 2 = 0 please show all work!

OpenStudy (gorv):

sin^2+cos^2=1 remember

OpenStudy (gorv):

cos^2=1-sin^2

OpenStudy (gorv):

now replace cos^2 2sin^2-(1-sin^2)-2=0 2sin^2-1+sin^2-2=0 3sin^2-3=0 3(sin^2-1)=0 sin^2-1=0 sin^2=1 sin=+-1 sin=n*pi/2

OpenStudy (gorv):

@ttop0816

OpenStudy (ttop0816):

thank you!